OpenStudy (anonymous):

A cyclist rides his bike at a rate of 22 feet per second. What is this rate in miles per hour? How many miles will the cyclist travel in 5 hours? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers.

OpenStudy (anonymous):

use conversion factors like this.\[\frac{ 22 feet }{ 1 second } \times \frac{ 1 mile }{ 5280 feet } \times \frac{ 3600 seconds }{ 1 hour }\] notice that all the units except mile and 1/hour "cancel out" leaving us with a final answer of mile/hour

OpenStudy (anonymous):

si u multiply al these?

OpenStudy (anonymous):

together?

OpenStudy (anonymous):

yes. when multiplying fractions this is equal to: \[\frac{ 22 \times 1 \times 3600 }{ 1 \times 5280 \times 1 }\]

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

@Euler271